OpenStudy (anonymous):

What is the greatest common factor of 18 and 36?

OpenStudy (cggurumanjunath):

18=2*3*3 36=2*2*3*3 so gcf is common terms in 18 and 36 they are :2*3*3=18

OpenStudy (anonymous):

so the greatest common factor is 18

OpenStudy (cggurumanjunath):

yes. :)

OpenStudy (anonymous):

wouldnt the greatest common factor be 9 since both 18 and 36 goes into 9??

OpenStudy (cggurumanjunath):

no

OpenStudy (anonymous):

why not??

OpenStudy (cggurumanjunath):

18 goes in 18 and 36

OpenStudy (cggurumanjunath):

see the first step !

OpenStudy (anonymous):

oh true blond moment.

OpenStudy (cggurumanjunath):

:)